2. Newsweek performed a poll in which 567 American parents were asked the question and found that only 12% agreed with the question , “Would you prefer to have your child taught by a male for grades K-12?”. Find the 95% C.I. for the true proportion of Americans agreeing with their child being taught by a male for grades K-12.
1
Expert's answer
2019-11-01T10:45:06-0400

Let "\hat{p}" be the sample proportion. When "n"  is large and "p" is not close to zero or one, we can use the normal distribution to approximate the binomial.

The sampling distribution of "\hat{p}" is normal.

The mean of the sampling distribution is "p."

The standard deviation of the sampling distribution (called the standard error) is


"SE=\sqrt{{p(1-p) \over n}}"

Given that "n=576, p=0.12."


"SE=\sqrt{{0.12(1-0.12) \over 576}}\approx0.013540"

Confidence Level 95%: "z^*-value=1.96"

Confidence interval for "p" is


"\hat{p}\pm z^*\times SE"

"0.12\pm 1.96\times 0.013540\approx0.12\pm0.0265"

"95\%\ CI=(0.0935, 0.1465)"